OpenStudy (anonymous):

integerate 3x^2/(1+x^6) dx

hartnn (hartnn):

did you try substituting, x^3 = u ??? du = ... ? x^6 = u^2

OpenStudy (anonymous):

du=3x^2

hartnn (hartnn):

du = 3x^2 dx

OpenStudy (anonymous):

therefore dx=du/3x^2

hartnn (hartnn):

go on, ask if you get stuck

OpenStudy (anonymous):

\[\int\limits (1+u^2)^-1 du\]

hartnn (hartnn):

yeah, integral 1/(1+u^2) dx thats the standard integral, do you know what this equals ?

hartnn (hartnn):

i mean du **

hartnn (hartnn):